OpenStudy (anonymous):

if a object is launched from a height of 125 meters horizontaly with a initial velocity of 90m/s....... ^x=??? ^t=???

OpenStudy (anonymous):

It will go as far as it can before it falls the 125 metres. s = u*t +1/2 * a * t^2 125 = 0 +1/2. 9.81 * t^2 SO now you have the time and can find out how far it has gone horizontally by using the same formula in that direction.